OpenStudy (anonymous):

What is the value of x for x^(2/3) = 3? Can anyone show me the steps to solve this problem please?

OpenStudy (jkristia):

Here is how I would solve this \[x^{\frac{2}{3}}=3\] Take log on both sides, bring down the exponent \[\frac{2}{3}\log(x)=\log(3)\] divide to isolate x \[\log(x) = \frac{\log(3)\times3}{2}\] raise both sides. 10 raised to log x is x. \[x = 10^{\frac{\log (3)\times3}{2}}\approx5.19615\] I'm not sure if there is an easier way to do this.

OpenStudy (anonymous):

Another way to do this is: Since \[(x^{m})^{n} = x^{mn}\] then \[(x^{2 \over 3})^{3 \over 2} = x^{{2 \over 3} \times {3 \over 2}} = x^1 = x\] So \[ x^{2 \over 3} = 3\] \[(x^{2 \over 3})^{3 \over 2} = 3^{3 \over 2}\] \[x = (3^3)^{1 \over 2}=\sqrt {27} \approx 5.196\]

OpenStudy (jkristia):

ahh - of course, and that is simpler.